OpenStudy (anonymous):

Is (2,-3) a solution of x+5y=4? **Medals Will Be Given**

OpenStudy (anonymous):

well to test this you will input x=2 y=-3 2+5(-3)=4 2-15=4 -9=4 this is false so (2,-3) is not a solution